0

ü c ŁóŚŁþ ý c ÃŁČŁõŊÙŁœď ďć ÝŁÑŁţŁĂŁõ ý c ŚŁû^{∗}öŚŁÞŁîSSŁČŁÚŁœď ćcŘÀŁãŁŮ ýŘÂŁŞ üŁĂŁČŁþŚŁŢ À cŁ§

K_{n} éc ŘÂŁð

öŚŁþćĄŁÞŁ¡Łõ ĆŁÜŁóŘćŚŁśŁä ÀŁśŁúŁÔŁĺŘ ćŘÀŁúŁŞ

ùÀČØİ

SSŁþŘ ďć ,ÀŁĂŁłŚŁŞüŁõ ŰŚŁČŁśŁČŁîÂŁŮ ýŚŁûÇŁ£ŁŞ SSŁþÂŁŮćÂŁŞďŚŁîÂŁŢ ø SSŁþÂŁŮřŁóŚŁŻ ěŘ üŁØŁþ ŚŁûöŊ ŚŁŞ ÍŁśŁŮÂŁõ ű c Ł§ŚŁśŁõ ø SSŁČŁŮĎ ýŚŁûâŁŞÂŁõ ďć öŚŁÞŁîSSŁČŁÚŁœď ÀŁĂŁİ SS cŁţŁêŚŁþ ýŘÂŁŞ üŁłøď ÅŁşŁĺ ,ÝŁþĄŁłüŁõ ŚŁĂŁłŊ öŚŁÞŁîSSŁČŁÚŁœď ø üŁţŁäŚŁĺ ťc ŁĂŁÞŁœďĄŁŮ ý c ÀŁĂŁŞöŚŁõě Ó c ŁþďŚŁãŁŮ ŚŁŞ ĆŁóŚŁÖŁõ

.ÝŁČŁûćüŁõ ĆŁþŘďŘ K_{n} éŘÂŁð

öŚŁÞŁîSSŁČŁÚŁœď ø üŁţŁäŚŁĺ ťc ŁĂŁÞŁœďĄŁŮ ý c ÀŁĂŁŞöŚŁõě 1

ćĄŁŁŻø ýďĄŁŁúŁŁÈŁŁõ ýĆŁŁČŁŁÌŁŁì .ÛŁŁõŚŁŁî Õ cŁŁŞŚŁŁÎŁŁŮ n − 1 ĆŁŁŞ K_{n} ýŚŁŁûñŚŁŁþ ěcŘÂŁŁêŘ ŚŁŁŞ ťŁŁĺŘ ÂŁŁŞŘÂŁŁŞ ťŁŁĂŁŁÞŁŁœďĄŁŁŮ ŒŁŁþ ýŘÂŁŁŞ ýÀŁŁĂŁŁŞöŚŁŁõě ŒŁŁþ SSŁŁŁþŘ ěŘ ĆŁŁŁœĄŁŁŁÞŁŁŁœ ŒŁŁŁþ ťc ŁŁŁąŚŁŁŁĺ ľc øď .ťŁŁŁĺŘ ÙŁŁŁœď n − 1 ŚŁŁŁŞ ýÃŁŁŁČŁŁŁõŊÙŁŁŁœď Û cŁŁŁŞŚŁŁŁì K_{n} ýŚŁŁŁûñŚŁŁŁþ ,Žøě n > 4 ÂŁŁŁû ýŘÂŁŁŁŞ ĆŁŁŁî ćďŘć

:ťŁĺŘ ùÀŁõŊ ÂŁþě Ý cŁó cďć (ýÀŁĂŁŞöŚŁõě)ýÃŁČŁõŊÙŁœď

.ťŁĺŘ ÙŁœď n − 1 ŚŁŞ ýÃŁČŁõŊÙŁœď Û cŁŞŚŁì K_{n} ýŚŁûñŚŁþ ,Žøě n > 4 ÂŁû ýŘěŘ ĆŁŞ .1 ÝŁó

K_{n} ěŘ ýÃŁČŁõŊÙŁœď ŒŁþ .ÝŁČŁĂŁîüŁõ ýďŘÁŁðùďŚŁÞŁł {∞, 0, 1, 2, …, n − 2} ýĆŁäĄŁÞŁźŁõ Â cŁňŚŁĂŁä ŚŁŞ Řď K_{n} ýŚŁûğŘď .ŰŚŁśŁŹŘ ěŘ ÀŁĂŁŮďŚŁśŁä ,(0 ≤ i ≤ n − 2) i Ù c Łœď ŚŁŞ ýŚŁûñŚŁþ :ÝŁČŁĂŁîüŁõ üŁêÂŁãŁõ ŰďĄŁň SSŁþŘ ĆŁŞ

\{\{∞, i\}, \{i − 1, i + 1\}, \{i − 2, i + 2\}, …, \{i − \frac{n}{2} + 1, i + \frac{n}{2} − 1\}\} (mod n − 1)

Řď n − 2 ŚŁŮ 0 ýŚŁûğŘď ø ÝŁČŁûć ďŘÂŁì ùÂŁþŘć ŒŁþ ÃŁîÂŁõ Řď ∞ ğ c Řď ÂŁðŘ ø ÝŁČŁþĄŁðüŁõ üŁţŁäŚŁĺ ý c ÃŁČŁõŊÙŁœď ,ýÃŁČŁõŊÙŁœď SSŁþŘ ĆŁŞ ĆŁŁČŁŁÜŁŁî ø ÀŁŁĂŁŁîüŁŁõ ÛŁŁňø i ğ c Řď ĆŁŁŞ Řď ÃŁŁîÂŁŁõ ĆŁŁî üŁŁóŚŁŁþ ŚŁŁŞ ÀŁŁœÂŁŁŞŘÂŁŁŞ i ÙŁŁœď ŚŁŁŞ ýŚŁŁûñŚŁŁþ ,ÝŁŁþďŘÁŁŁÚŁŁŞ ùÂŁŁþŘć Íc ŁŁČŁŁ¡ŁŁõ ýøď ÂŁŁŞŘÂŁŁŞ Û cŁŁňŘĄŁŁê ŚŁŁŞ

:ĆŁî ćĄŁŞ ÀŁĂŁûŘĄŁą ÙŁœÂŁÞŁû üŁŮďĄŁň ďć ñŚŁþ øć ťŁĺŘ ćÂŁê n − 1 öĄŁİ ĆŁźŁČŁţŁœ ďć .öŊ ÂŁŞ ćĄŁÞŁä ýŚŁûñŚŁþ ýÃŁŁîÂŁŁõÂŁŁČŁŁè ñ cŚŁŁþ ýÂŁŁÚŁŁþć ø (ťŁŁĺŘ ∞ öŊ ÂŁŁĺ ŒŁŁþ)ýÃŁŁîÂŁŁõ ñ cŚŁŁþ ŚŁŁûöŊ ěŘ üŁŁØŁŁþ ŰďĄŁŁň SSŁŁþŘďć ĆŁŁî ,ÀŁŁĂŁŁłŚŁŁŞ ćĄŁŁÞŁŁä ÝŁŁû ÂŁŁŞ (1

Transversal^{∗}

.ťŁĺŘ

.ÀŁĂŁţŁÆŁû ýÃŁîÂŁõÂŁČŁè ñ cŚŁþ øć ÂŁû ŰďĄŁň SSŁþŘ ďć ĆŁî ,ÀŁĂŁłŚŁŞ ýěŘĄŁõ ÝŁû ŚŁŞ (2

ğ c ĄŁŁŐď ćŘÀŁŁãŁŁŮ Û cŁŁìŘÀŁŁ§ ýĆŁŁêŚŁŁőŘĆŁŁŞ 1 ÂŁŁŞŘÂŁŁŞ Řď ýÃŁŁîÂŁŁõÂŁŁČŁŁè ñ cŚŁŁþ ŒŁŁþ ñ cĄŁŁÏ ÂŁŁðŘ ø ÀŁŁĂŁŁţŁŁÆŁŁČŁŁœ ÙŁŁœÂŁŁÞŁŁû ýÃŁŁîÂŁŁõ ñ cŚŁŁþ øć žŁŁČŁŁû SSŁŁČŁŁĂŁŁżŁŁÞŁŁû ďć ÂŁŁÔŁŁň Řď ýÃŁŁîÂŁŁõ ýŚŁŁûñŚŁŁþ ñ cĄŁŁÏ ø ÀŁŁĂŁŁîüŁŁõ ÛŁŁňø ÂŁŁÚŁŁþć Â cŁŁĺ ĆŁŁŞ Řď ñŚŁŁþ Â cŁŁĺ ŒŁŁþ ùÂŁŁþŘć Íc ŁŁČŁŁ¡ŁŁõ ýøď ĆŁŁî ÝŁŁþÂŁŁČŁŁÚŁŁŞ ýÂŁŁČŁŁÆŁŁõ ýøď ŚŁŞ ñ cŚŁþ øć žŁČŁû ĆŁźŁČŁţŁœ ďć ø ,ćďŘć ĄŁÌŁä ŒŁþ ŚfŁÖŁČŁìć ôŘÀŁî ÂŁû ěŘ ,\frac{n}{2} − 1 ŚŁŮ 0 ñĄŁÏ ĆŁŞ ýŚŁûñŚŁþ ěŘ ÙŁœď ÂŁû ùŚŁÚŁœŊ ,ÝŁþÂŁČŁÚŁŞ ÂŁÑŁœ

.ÀŁĂŁţŁÆŁČŁœ ÙŁœÂŁÞŁû öŚŁĺŒŁþ ñ cĄŁÏ

[figure: ∞; 0; 1; 2; 3; 4; K_{6} ďć 0 ÙŁœď ýŚŁúŁóŚŁþ]

□ .ÀŁœďŘÀŁœ üŁœŚŁĺŒŁþ Ù c Łœď öŊ ďć üŁóŚŁþ øć žŁČŁû ĆŁî ťŁĺŘ üŁÜŁõŚŁî Õ cŁŞŚŁÎŁŮ ,éŘÂŁð ŒŁþ üŁóŚŁþ ýÃŁČŁõŊÙŁœď ŒŁþ ďć öŚŁÞŁîSSŁČŁÚŁœď ŒŁþ .ÀŁŁĂŁŁîüŁŁõ ěŘÂŁŁêŘ Řď éŘÂŁŁð ýŚŁŁûñŚŁŁþ ĆŁŁî ťŁŁĺŘ n − 1 ,K_{n} éŘÂŁŁð ďć ŘÃŁŁźŁŁõñŚŁŁþ øćĆŁŁŞøć ýŚŁŁûöŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď SSŁŁØŁŁÞŁŁõ ćŘÀŁŁãŁŁŮ ÂŁŁůŁŁîŘÀŁŁ§ ďć ŘÃŁźŁõñŚŁþ ö cŚŁÞŁîSSŁČŁÚŁœď ÀŁĂŁİ ćcĄŁŻø .ćďŘć öŚŁÞŁîSSŁČŁÚŁœď ŒŁþ ÛŁìŘÀŁ§ K_{n} üŁóŚŁþ ýÃŁČŁõŊÙŁœď ÂŁû ĆŁî ťŁĺŘ ùÀŁł ťŁŞŚŁŹ SSŁČŁĂŁżŁÞŁû ŘÃŁźŁõñŚŁþ øćĆŁŞøć ö cŚŁÞŁîSSŁČŁÚŁœď ÀŁĂŁİ K_{n} ĆŁî ťŁĺŘ ùÀŁÈŁœ ÉŁ£ŁÈŁõ ěĄŁĂŁû âŁìŘø ďć .ťŁĺŘ ŚŁõ ű c Ł¡ŁŞ áĄŁőĄŁõ üŁţŁäŚŁĺ ýÃŁČŁõŊÙŁœď

.ťŁĺŘ ŘÃŁźŁõñŚŁþ øćĆŁŞøć ýŚŁûöŚŁÞŁîSSŁČŁÚŁœď ĆŁŞ ěŘÂŁêŘ ÛŁŞŚŁì K_{n} üŁþŚŁûn ĆŁİ ýŘÂŁŞ ŚŁþ ø ćďŘć

K_{n} üŁţŁäŚŁĺ ýÃŁČŁõŊ ÙŁœď ďć ÝŁÑŁţŁĂŁõ ýŚŁûöŚŁÞŁîSSŁČŁÚŁœď 2

ÛŁČŁØŁÈŁŮ ÂŁŞŘÂŁŞ ñ cĄŁÏ ŚŁŞ ñŚŁþ \frac{n}{2} − 1 ø 0 ñ cĄŁÏ ĆŁŞ ñŚŁþ ŒŁþ ěŘ ĆŁî üŁœŚŁÞŁîSSŁČŁÚŁœď :ÝŁČŁĂŁîüŁõ ÓŁþÂŁãŁŮ SSŁČŁĂŁİ Řď ÝŁÑŁţŁĂŁõ ö cŚŁÞŁîSSŁČŁÚŁœď ŒŁþ ćcďĄŁõ ďć üŁóø ÀŁłŚŁŞ ŘÃŁźŁõñŚŁþ ö cŚŁÞŁîSSŁČŁÚŁœď ÀŁĂŁİ ĆŁŞ ěŘÂŁêŘ Û cŁŞŚŁì K_{n} ĆŁî ýŘn > 2 ÀŁłŚŁŞ ĆŁţŁłŘÀŁœ ćĄŁŻø ÀŁĺďüŁõ ÂŁÑŁœ ĆŁŞ .ÀŁłŚŁŞ ùÀŁł

:ćĄŁłüŁõ ŰŚŁśŁŹŘ üŁţŁ§Řď ĆŁŞ áĄŁőĄŁõ SSŁþŘ ÝŁÑŁţŁĂŁõ ýŚŁûöŚŁÞŁîSSŁČŁÚŁœď

.ťŁÆŁČŁœ ÝŁÑŁţŁĂŁõ ýŚŁûöŚŁÞŁîSSŁČŁÚŁœď ĆŁŞ ěŘÂŁêŘ Û cŁŞŚŁì K_{n} ,ýŘn > 2 žŁČŁû ýŘěŘ ĆŁŞ .1 ĆŁČŁÌŁì

,ÀŁœøď ďŚŁîĆŁŞ ÀŁþŚŁŞ ÝŁû k ñ cĄŁÏ ĆŁŞ ýŚŁûñŚŁþ ýĆŁČŁÜŁî ø ćďŘć (k)ÂŁŞŘÂŁŞ ñ cĄŁÏ ŚŁŞ ñŚŁþ \frac{n}{2} − 1 ÝŁÑŁţŁĂŁõ ö cŚŁÞŁîSSŁČŁÚŁœď ÂŁû ÝŁČŁœŘćüŁõ .ŰŚŁśŁŹŘ ćcŘÀŁŁãŁŁŮ SSŁŁþÂŁŁţŁŁÈŁŁČŁŁŞ)ťŁŁóŚŁŁ§ SSŁŁþÂŁŁţŁŁúŁŁŞ ďć âŁŁìŘø ďć ,ťŁŁÆŁŁČŁŁœ ÂŁŁþÁŁŁŢÇŁŁ£ŁŁŞ \frac{n}{2} − 1 ÂŁŁŞ ĆŁŁî ťŁŁĺŘ n − 1 ,k ñĄŁŁÏ ĆŁŁŞ ýŚŁŁûñŚŁŁþ ćŘÀŁŁãŁŁŮ üŁŁóø ĆŁŞ ŰďĄŁň SSŁþŘ ďć ĆŁî ÝŁČŁłŚŁŞ ĆŁţŁłŘć ŘÃŁźŁõñŚŁþ Ý cŁÑŁţŁĂŁõ ö cŚŁÞŁîSSŁČŁÚŁœď n − 2 ÝŁČŁœŘĄŁŮüŁõ (SSŁØŁÞŁõ ýŘÃŁźŁõñŚŁþ Ý cŁÑŁţŁĂŁõ ýŚŁûöŚŁÞŁîSSŁČŁÚŁœď ö cŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď (n − 2)ćŘÀŁŁãŁŁŮ SSŁŁþŘ .ćďŘÀŁŁœ ťŁŁîÂŁŁł üŁŁœŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď žŁŁČŁŁû ďć k ñ cĄŁŁÏ ĆŁŁŞ ñŚŁŁþ ŒŁŁþ ŚfŁŁÖŁŁČŁŁìć 0 ≤ k ≤ \frac{n}{2} − 1 ÂŁŁû ýŘěŘ □ .ćďŘć ćĄŁŻø n = 6 ýŘÂŁŞ ŘÃŁźŁõñŚŁþ Ý cŁÑŁţŁĂŁõ k ñ cĄŁŁŁÏ ĆŁŁŁŞ ñŚŁŁŁþ \frac{n}{2} − 1 ø \{0, ∞\} \begin{matrix}ñ \\ c \end{matrix}ŚŁŁŁþ ěŘ ĆŁŁŁî K_{n} ďć üŁŁŁÜŁŁŁõŚŁŁŁî ÕŁŁŁŞŚŁŁŁÎŁŁŁŮ ŚŁŁŁŞ ťŁŁŁĺŘ ÂŁŁŁŞŘÂŁŁŁŞ (2k|n − 2 ø Žøě n) T_{n,k} .ÓŁŁŁþÂŁŁŁãŁŁŁŮ

:ťŁĺŘ ÂŁþě Û cŁØŁł ĆŁŞ ø ùÀŁł ÛŁČŁØŁÈŁŮ

\{\{0, ∞\}\} + \{\{i + 2jk, i + (2j + 1)k\}|1 ≤ i ≤ k, i + (2j + 1)k ≤ n − 2, 0 ≤ j ≤ \frac{n − 2}{2k} \}

.ťŁĺŘ öŚŁÞŁîSSŁČŁÚŁœď ùďŘĄŁÞŁû T_{n,k} ,n = 4m + 2 ÂŁðŘ üŁóø ťŁÆŁČŁœ öŚŁÞŁîSSŁČŁÚŁœď ŒŁþ ùŚŁðžŁČŁû T_{n,k} ,n = 4m ÂŁðŘ .2 ĆŁČŁÌŁì

ÀŁŁłŚŁŁŞ α ŚŁŁûĆŁŁţŁŁĺć ćŘÀŁŁãŁŁŮ ÝŁŁČŁŁĂŁŁî üŁŁõ ŋÂŁŁê .ťŁŁĺŘ ùÀŁŁł ùćŘć öŚŁŁÈŁŁœ T_{n,k} ďć ĆŁŁţŁŁĺć ýŚŁŁĂŁŁãŁŁõ ÂŁŁþě Û cŁŁØŁŁł Œ c ŁŁþ ťc ŁŁÞŁŁÆŁŁì ďć .ŰŚŁŁśŁŁŹŘ T_{n,k} ÅŁŁŢ ÀŁŁĂŁŁłŚŁŁŞ ÙŁŁœÂŁŁÞŁŁû ÀŁŁĂŁŁœŘĄŁŁŮüŁŁÞŁŁœ k ñ cĄŁŁÏ ĆŁŁŞ ýŚŁŁûñŚŁŁþ ĆŁŁî ŚŁŁźŁŁœŊ ěŘ .ćďŘć ćĄŁŁŻø ğŘď 2k ĆŁŁţŁŁĺć ÂŁŁû ďć ø α = \frac{n−2}{2k} ÅŁŁŢ SSŁŁČŁŁŞ üŁŁóŚŁŁþ žŁŁČŁŁû öĄŁŁİ .ÀŁŁłŚŁŁśŁŁœ ÙŁŁœÂŁŁÞŁŁû k ñ cĄŁŁÏ ĆŁŁŞ ýŚŁŁûñŚŁŁþ ěŘ ôŘÀŁŁî žŁŁČŁŁû ŚŁŁŞ \{0, ∞\} \begin{matrix}ñ \\ c \end{matrix}ŚŁŁþ ÂŁŁðŘ ÍŁŁÖŁŁê ø ÂŁŁðŘ ťŁŁĺŘ öŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď ñŚŁþ ÂŁŞ ćĄŁÞŁä üŁóŚŁþ üŁţŁìø ŚŁúŁĂŁŮ ÀŁþć öŘĄŁŮüŁõ üŁţŁ§Řď ĆŁŞ .ÝŁČŁĂŁîüŁõ üŁĺďÂŁŞ Řď ĆŁţŁĺć ÂŁû öøďć ýŚŁûñŚŁþ ćďŘÀŁœ ćĄŁŻø ĆŁţŁĺć øć ğĄŁŐď n − 2 ŚŁŁŮ \frac{n−2}{2} + 1 ø \frac{n−2}{2} ŚŁŁŮ 1 ğĄŁŁŐď SSŁŁČŁŁŞ üŁŁóŚŁŁþ ÀŁŁłŚŁŁŞ Žøě α ÂŁŁðŘ öĄŁŁİ ÀŁŁĂŁŁłŚŁŁŞ ćÂŁŁê øć ÂŁŁû k ø α ĆŁŁî ćďŘć ćĄŁŁŻø \{0, ∞\}

.ćĄŁłüŁÞŁœ {0, ∞} ýěŘĄŁõ üŁóŚŁþ ÝŁû ÀŁłŚŁŞ Žøě k ø ćÂŁê α ÂŁðŘ ø ćďŘÀŁœ ćĄŁŻø

üŁóŚŁþ ÅŁŢ ÀŁĂŁţŁÆŁû ćÂŁê α ø k ĆŁźŁČŁţŁœ ďć ,kα = 2m − 1 ÅŁŢ 2kα = n − 2 = 4m − 2 ŰďĄŁň SSŁþŘ ďć :n = 4m (1 .ťŁÆŁČŁœ öŚŁÞŁîSSŁČŁÚŁœď ŒŁþ T_{n,k} ćĄŁłüŁõ ĆŁźŁČŁţŁœ ø ćďŘć ćĄŁŻø {0, ∞} ŚŁŞ ÙŁœÂŁÞŁû

ťŁĺŘ Žøě α ø k ěŘ üŁØŁþ ÛŁìŘÀŁ§ ĆŁźŁČŁţŁœ ďć ,kα = 2m ÅŁŢ 2kα = n − 2 = 4m ŰďĄŁň SSŁþŘ ďć :n = 4m + 2 (2 .ťŁĺŘ öŚŁÞŁîSSŁČŁÚŁœď ŒŁþ T_{n,k} ćĄŁłüŁõ ĆŁźŁČŁţŁœ ø ćďŘÀŁœ ćĄŁŻø {0, ∞} ŚŁŞ ÙŁœÂŁÞŁû üŁóŚŁþ ÅŁŢ

[figure: ćÂŁê k ,ćÂŁê α (4; Žøě k ,ćÂŁê α (3; Žøě α (2; ĆŁţŁĺć ÓŁþÂŁãŁŮ (1; ñøaŘ ýĆŁţŁĺć; ÂŁąŊ ýĆŁţŁĺć; (ôŘe\frac{n−2}{2k})]

□ ÀŁþć öŘĄŁŮüŁõ öďŚŁÖŁŮ ťŁÜŁä ĆŁŞ .ÀŁœŘùÀŁł âŁÞŁŻ i ŚŁŞ (n − 1) ÙŁĂŁû ĆŁŞ üŁÚŁÞŁû öŊ ćŘÀŁäŘ ĆŁî ťŁĺŘ T_{n,k} öŚŁÞŁû :T_{n,k}^{(i)} .ÓŁþÂŁãŁŮ

.ÀŁłŚŁŞ öŚŁÞŁîSSŁČŁÚŁœď T_{n,k} ÂŁðŘ ÍŁÖŁê ø ÂŁðŘ ťŁĺŘ öŚŁÞŁîSSŁČŁÚŁœď T_{n,k}^{(i)} ,i ∈ Z ÂŁû ýŘěŘ ĆŁŞ ĆŁî

ø ùćĄŁŞ öŚŁÞŁîSSŁČŁÚŁœď ŒŁþ üŁþŚŁúŁĂŁŮ ĆŁŞ ôŘÀŁîÂŁû ÂŁðŘ ÍŁÖŁê ø ÂŁðŘ ÀŁĂŁţŁÆŁû ŘÃŁźŁõñŚŁþ ö cŚŁÞŁîSSŁČŁÚŁœď øć T_{n,k′}^{(i′)} ø T_{n,k}^{(i)} .2 ÝŁó

i ≠ i′ ùŚŁÚŁœŊ k ≠ k′ (1
|i − i′| ≡ k (mod n − 1) ùŚŁÚŁœŊ k = k′ (2

SSŁþŘ ,i ≠ i′ ÂŁðŘ ĆŁî ÀŁĂŁţŁÆŁû {∞, i′} ø {∞, i} ýŚŁûñŚŁþ öŚŁÞŁîSSŁČŁÚŁœď øć ñ cĄŁÏÝŁû ýŚŁûñŚŁþ ŚŁúŁĂŁŮ ĆŁźŁČŁţŁœ ďć k ≠ k′ ÂŁðŘ .ŰŚŁśŁŹŘ ñŚŁŁþ ŒŁŁþ ďć ÛŁŁìŘÀŁŁ§ öŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď øć ùŚŁŁðöŊ |i − i′| ≢ k (mod n − 1) ÂŁŁðŘ ,k = k′ ÂŁŁðŘ üŁŁóø .ćĄŁŁŞ ÀŁŁĂŁŁûŘĄŁŁą ŘÃŁŁźŁŁõ ÃŁŁČŁŁœ øć □ .ćĄŁŞ ÀŁĂŁûŘĄŁą ŘÃŁźŁõñŚŁþ ,øć SSŁþŘ ĆŁî ÀŁþć öŘĄŁŮüŁõ T_{n,k} Ó c ŁþÂŁãŁŮ ÕŁśŁÏ ŰďĄŁňSSŁþŘ ÂŁČŁè ďć üŁóø ťŁłŘć ÀŁĂŁûŘĄŁą íŘÂŁţŁłŘ k ñ cĄŁÏ ĆŁŞ

ö cŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď \frac{n−2}{2} ýŚŁŁûĆŁŁČŁŁÜŁŁäôĄŁŁÆŁŁÖŁŁõ ćŘÀŁŁãŁŁŮ ýùøĚŁŁäĆŁŁŞ 4 ÛŁŁìŘÀŁŁ§ K_{n} ,n > 10 ĆŁŁî n = 4m + 2 ÂŁŁû ýŘěŘ ĆŁŁŞ .3 ĆŁČŁÌŁì .ćďŘć ŘÃŁźŁõñŚŁþ Ý cŁÑŁţŁĂŁõ

.ÀŁŁĂŁŁûćüŁŁõ ŘÃŁŁźŁŁõñŚŁŁþ ö cŚŁŁÞŁŁîSSŁŁČŁŁÚŁŁœď 8 Û cŁŁČŁŁØŁŁÈŁŁŮ T_{n, \frac{n−2}{4}}^{(5+ \frac{n−2}{4} )} ø T_{n, \frac{n−2}{4}}^{(5)} ,T_{n, \frac{n−2}{2}}^{(3+ \frac{n−2}{2} )} ,T_{n, \frac{n−2}{2}}^{(3)} ,T_{n,2}^{(4)} ,T_{n,2}^{(2)} ,T_{n,1}^{(1)} ,T_{n,1}^{(0)} .ŰŚŁŁśŁŁŹŘ ŘÀŁŁŁŁČŁŁŁŁŢ ùÀŁŁŁŁÈŁŁŁŁœ ùćŚŁŁŁŁÔŁŁŁŁţŁŁŁŁĺŘ \{∞, i\} \begin{matrix}ñ \\ c \end{matrix}ŚŁŁŁŁþ ĆŁŁŁŁî Řď i ŒŁŁŁŁþ ,\frac{n−2}{2} ø \frac{n−2}{4} ,2 ,1 ěŘ ÂŁŁŁŁČŁŁŁŁè ,k Û cŁŁŁŁůŁŁŁŁõ \frac{n−2}{2} Ć cŁŁŁŁČŁŁŁŁÜŁŁŁŁäôĄŁŁŁŁÆŁŁŁŁÖŁŁŁŁõ ÂŁŁŁŁû ýŘěŘ ĆŁŁŁŁŞ ñŚŁŁŁŁ§ ŠŚŁ£ŁţŁœŘ ŘÃŁźŁõñŚŁþ ö cŚŁÞŁîSSŁČŁÚŁœď n − 1 ĆŁî ťŁĺŘ üŁĂŁãŁõ SSŁþŘ ĆŁŞ ÀŁłŚŁŞ ĆŁţŁłŘÀŁœ ćĄŁŻø ÂŁðŘ ŘÂŁþě ćďŘć ćĄŁŻø ŚŁÞŁţŁ§ ýŘi SSŁČŁĂŁİ)ùćÂŁî □ .ťŁĺŘ ùÀŁł ĆŁţŁĺŘĄŁą ćcŘÀŁãŁŮ öŚŁÞŁû ,ćŘÀŁãŁŮ SSŁþŘ ,ÝŁČŁþŘÃŁêŘüŁõ öŚŁÞŁþŚŁûöŚŁÞŁîSSŁČŁÚŁœď ĆŁŞ Řď T_{n,K}^{(i)} ø (ÝŁþŘùćÂŁî

ýŘÃŁźŁõñŚŁþ øćĆŁŞøć ýŚŁûöŚŁÞŁîSSŁČŁÚŁœď ćcŘÀŁãŁŮ Â cŁůŁîŘÀŁ§ ñ cøÀŁŻ ŒŁİĄŁî ýŚŁûn ýŘÂŁŞ (t_{n}) K_{n} éŘÂŁð

\begin{matrix}t_{n} & n \\ 1 & 2 \\ 0 & 4 \\ 4 & 6 \\ 3 & 8 \\ ≥ 8 & 10 \\ ≥ 10 & 12 \\ ≥ 11 & 14 \\ ≥ 12 & 16 \\ ≥ 13 & 18 \\ ≥ 14 & 20 \\ ≥ 15 & 22 \\ ≥ 15 & 24 \\ ≥ 15 & 26 \end{matrix}

âŻŘÂõ

ñŚŁĺ üŁőŚŁþď üŁœŚŁúŁŻ ćcŚŁČŁşŁÞŁóŘ ýĆŁóŚŁÆŁõ ŒŁþ ýŚŁûĆŁÈŁþď ,öŚŁþÀŁúŁõ ÀŁÞaŁ¡Łõ ğÀŁĂŁúŁõ ø öŚŁþćĄŁÞŁ¡Łõ ĆŁóŘćŚŁśŁä ÀŁČaŁĺ ÂŁţŁîć ]1[ ,ùøÄŁŁŢÇŁŁœŘć ýĆŁŁõŚŁŁĂŁŁÜŁŁÊŁŁê ø 1377 ď cŚŁŁúŁŁŞ ,51 ýùďŚŁŁÞŁŁł ,ÝŁŁûćěŘøć ñ cŚŁŁĺ ,üŁŁőŚŁŁþď ľc ěĄŁŁõŊ À cŁŁłď ýĆŁŁÜŁŁźŁŁõ ,1997

.1377 ö cŚŁţŁÆŁõě ,2 ýùďŚŁÞŁł

[2] Charles J. Colbourn and Jeffrey H. Dinitz, Latin Squares, MOLs and Orthogonal arrays, The CRC Handbook of Combinatorial Designs.